OpenStudy (anonymous):

What is the limit as n approaches infinity of (3sqrt(n))^(1/2n) Main reason this frustrates me is cause I know a bunch of easy ways to do this I'm not allowed to use. I cannot use L'hopitals or even ln(f(x)). I have the epsilon definition of a limit and the basic addition/subtraction/multiplication/division limit laws

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} (3\sqrt{n})^{1/2n}\] is how it appears in my book

OpenStudy (anonymous):

Is it \(3\sqrt n\) or \(\sqrt[3]{n}\) ? Also, is it \(\dfrac{1}{2n}\) or \(\dfrac{1}{2}n\) ?

OpenStudy (anonymous):

I know its \[3\sqrt{n}\] I believe its \[\frac{ 1 }{ 2n }\] as otherwise the limit is divergent (+infinity) and the question is phrased "Determine the following limits". Forget to mention I do have squeeze theorem

OpenStudy (anonymous):

Those instructions don't seem to suggest that the limit must exist or be finite

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} (3\sqrt{n})^{1/2n}\] is how it appears in my book.

OpenStudy (anonymous):

I keep getting it to the point where I just need to prove that \[\lim_{n \rightarrow \infty} (9n)^{(1/n)}=1\] but I can't get there.without using one of the methods I'm not allowed to.

OpenStudy (anonymous):

So you have to both find the limit and prove it?

OpenStudy (anonymous):

Well I have to prove that the limit I found is correct, but that is usually proven in the work finding the limit. I may not be understanding you here.

OpenStudy (anonymous):

Allow me to clarify: evaluating the limit is a process completely distinct from proving a limit. Evaluation is done with algebraic manipulation until you have something that can be evaluated directly (provided the expression is continuous), while a proof would involve applying the \(\epsilon-\delta\) definition to establish that \(\lim_{n\to\infty}f(n)=1\).

OpenStudy (anonymous):

Ah, I can't pull 1 out of a straw hat unfortunately so we need to evaluate it. If at some point it becomes a rather simple limit where it is foreseeable without using the before mentioned out-lawed methods I may use the epsilon definition to confirm it.

OpenStudy (anonymous):

I hope that makes sense.

OpenStudy (anonymous):

You're not given a lot of freedom here... L'Hopital's would have been of great use. The algebraic properties of limits won't be much help. The squeeze theorem would probably be the most useful way to approach this, but it'll be tricky to come up with bounding sequences.

OpenStudy (anonymous):

Aye, I want to use the sequence (1) as the lower bound but an upper bound sequence that goes to 1....is rather difficult.

OpenStudy (anonymous):

I suppose there's nothing stopping you from using \(a_n=1-\dfrac{1}{n}\) and \(b_n=1+\dfrac{1}{n}\) (since we know that the limit must be 1.

OpenStudy (anonymous):

Or \(a_n=1\) and \(b_n=1+\dfrac{1}{n}\).

OpenStudy (anonymous):

alright, give me a sec to try and work on a proof that \[b _{n} \ge x _{n}\] for sufficiently large n

OpenStudy (anonymous):

You would then only need to show that for all \(n\), or at least all \(n\) after some large enough \(N\), the following holds. \[\large(3\sqrt n)^{\frac{1}{2n}}\le1+\frac{1}{n}\]

OpenStudy (anonymous):

I might suggest a proof by induction to establish this, but that may be easier said than done.

OpenStudy (anonymous):

definitely easier said then done, might try it on the equation \[(9n)^{1/n}\] as I do have a theorem that if \[\lim_{n \rightarrow \infty} x _n = c \rightarrow \lim_{n \rightarrow \infty} \sqrt x _n = \sqrt c\] there are some other conditions but this sequence satisfies them